OpenStudy (anonymous):

For what values of y does y^2-2y=48? Show all work.

OpenStudy (amistre64):

-6 and 8

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

he is super fast that is how. plus wolfram

OpenStudy (anonymous):

\[y^2-2y-48=0\] \[(y+6)(y-8)=0\]

OpenStudy (amistre64):

sometimes I feel like working it out; others times not lol

OpenStudy (anonymous):

\[y+6=0\] \[y=-6\] \[y-8=0\] \[y=8\]

OpenStudy (anonymous):

method two \[y^2-2y=48\] \[(y-1)^2=48+1=49\] \[y-1=7\] or \[y-1=-7\] \[y = 8\] or \[y=-6\]

OpenStudy (anonymous):

pfft super fast , anyone thats been doing maths for like 4yrs should be able to solve a quadratic like that in about 30seconds

OpenStudy (anonymous):

its even monic which makes it even easier!

OpenStudy (amistre64):

28.6875378583032847459504348475... secs lol

OpenStudy (amistre64):

30 seconds is for my grandma ;)

OpenStudy (anonymous):

i still cannot factor to save my life. prefer to find the zeros

OpenStudy (amistre64):

the factors are just the zeros in reverse; if x = -3, 5 are roots then (x+3)(x-5) are the factors

OpenStudy (amistre64):

maybe should ammend that to have a constant of variation, but you know lol

OpenStudy (anonymous):

of course and so to factor i find the zeros. it is easier for me to do it that way

OpenStudy (anonymous):

Satelite, thanks for the tip about wolfram. It calculated the answer to "life, the universe, and everything" super-fast !